Allissa:

The sum of two numbers is 28. The first number, x, is three times the second number, y. Which system of equations can be used to find the two numbers?

Symphony:

The sum of two numbers is 28. Lets assume the two unknown number as x and y sum means add when 2 unknown numbers is add is gives a value 28(according to the statement) Lets write it in eqn form \(x+y=28\) The first number, x, is three times the second number, y. okay,in order y to be equal to x, y needs to be multiplied three times the value. \(x=3y\)

Symphony:

x-3y=0

Symphony:

Therefore,the system of eqns can be used to find the two numbers are \(x+y=28\) \(x-3y=0\)

Symphony:

Hope this helps ^

jhonyy9:

nice work @Symphony ,understandably sure easy ... @Allissa hope this help you